One mole of a mixture of N2,NO2, & N2O has a mean molar mass of 55.4%.On heating to a temperature at which all the N2O4 may be pressumed to have dissiciated:N2O4 = 2NO2,the mean molar mass tends to the lower value of 39.6.What is the ratio of N2 : NO2 : N2O4 in the original mixture ?

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Mean molar mass = > n1(28) + n2(46) + n3(92)/(n1 +n2 +n3)  =  >  55.4..................(1)
and from N2O4 = 2NO2
               n3            n2 = 2*n3..................(2)
and n1 + n2 + n3 = 1.........................(3)
solving eqns. (1) (2) and (3) we get,
n1 = 0.19
n2 = 0.536
n3 = 0.274
the fratio would be=>
1 : 2.82 : 1.44